OpenStudy (anonymous):

Use f(x)=x+2 and g(x)=x+3 to find each of the following: a.(f+g)(x),b. (f-g)(x).c. (fg)(x),d. (f/g)(x)

OpenStudy (anonymous):

can someone help please

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

a. (f+g): so add the two functions: (x+2) +(x+3) =x+2+x+3= 2x+5

OpenStudy (anonymous):

b. f-g: subtract the two functions (x+2) - (x+3)= x+2 -x -3 = -1

OpenStudy (anonymous):

c. multiply the functions d. divide the functions if need more clarification just ask!

OpenStudy (anonymous):

i do need more clarification

OpenStudy (anonymous):

c. 6 and d is 1.5

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok I need more clarification

OpenStudy (anonymous):

c. (x+2)(x+3)= so u have to factor

OpenStudy (anonymous):